Consider 40 students in a Chemistry class that each have random amounts of liquid in their beakers. Suppose that the amount of liquid in each beaker is Normally distributed with mean 0.20 liters and standard deviation 0.05 liters.

a. Find the probability that the class has between 7.9 and 8.1 liters of liquid altogether.

b. Find the value of b such that there is a 95 percent chance that total quantity of liquid is between 8?b and 8+b liters altogether.
Answer #1

Ans:

mean for sum of all liquid amounts=40*0.2=8

standard deviation=40*0.05=2

a)

z(7.9)=(7.9-8)/2=-0.05

z(8.1)=(8.1-8)/2=0.05

P(-0.05<z<0.05)=P(z<0.05)-P(z<-0.05)=0.5199-0.4801=0.0399

b)

95% of the data falls within 2 standard deviations of the mean.

standard deviation=2

lower limit=8-2*2=4

upper limit=8+2*2=12

95% 95 percent chance that total quantity of liquid is between 4 and 12 liters altogether.